# Integrating Multimodal Data for Joint Generative Modeling of Complex Dynamics

### v2

@@ A.1 MTF framework
The MVAE consists of a set of decoder models, $p_{\bm{\theta}}(\bm{Y}|\bm{\tilde{Z}})$, shared with the DSR model, with examples given in Eqs. 5-7 and above, a prior model $p_{\bm{\theta}}(\bm{\tilde{Z}})$ instantiated by the DSR model, see eq. 9, and the encoder (approximate posterior) $q_{\bm{\phi}}(\bm{\tilde{Z}}|\bm{Y})$, where a sample from this distribution provides the TF signal (see also Fig. 1). This setup is modular and all model components may be flexibly chosen. We assume conditional independence (given latent states) of the observations, such that the likelihood terms related to the observations simply sum up, i.e. $\log p_{\bm{\theta}}(\bm{{Y}}|\bm{\tilde{Z}})=\sum_{t=1}^{T}\left(\log p_{\bm{\theta}}(\bm{x}_{t}|\bm{\tilde{z}}_{t})+\log p_{\bm{\theta}}(\bm{o}_{t}|\bm{\tilde{z}}_{t})+\log p_{\bm{\theta}}(\bm{{c}_{t}}|\bm{\tilde{z}}_{t})\right)$ for the decoders specified in Eqs. 5-7 (with parameters shared with Eqs. 2-4).
For the encoder part, as the latent dendPLRNN process itself is conditionally Gaussian, we make a Gaussian assumption for the variational density $q_{\bm{\phi}}(\bm{\tilde{Z}}|\bm{Y})=\mathcal{N}(\bm{\mu}_{\bm{\phi}}(\bm{Y}),\bm{\Sigma}_{\bm{\phi}}(\bm{Y}))$ to approximate the true posterior $p(\bm{\tilde{Z}}|\bm{Y})$, where mean and covariance are functions of the data. We further use the common mean field approximation to factorize $q_{\bm{\phi}}(\bm{\tilde{Z}}|\bm{Y})$ across time (Girin et al., 2021). The mean $\bm{\mu}_{\theta}(\bm{Y})$ and covariance $\bm{\Sigma}_{\theta}(\bm{Y})$ of the approximate posterior are parameterized by one-dimensional (across time) convolutional layers (Brenner et al., 2022), using observations within the window defined by the input kernel. For the mean we use a 4-layer stack of convolutional layers with decreasing kernel sizes $11,8,5,3$, while the diagonal covariance was parameterized by a single convolutional layer with a kernel size of $11$. We also tested various other choices for the encoder, as discussed in Suppl. A.4.
@@
For the initial condition $\bm{z}_{1}$ of the latent process, if $K<M$, the $M-K$ remaining states are randomly sampled from a standard normal distribution. As shown in Mikhaeil et al. (2022), the best tradeoff between exploding gradients and capturing relevant long-term dependencies is achieved when choosing the forcing interval $\tau$ according to the system’s maximal Lyapunov exponent (predictability time). However, for non-continuous data (like counts), this cannot readily be determined, in which case the forcing interval $\tau$ may be regarded as a hyper-parameter determined via grid search (see also Suppl. Fig. A5 for example reconstructions for different choices of $\tau$).
@@ A.2 Performance measures
The power spectrum Hellinger distance ($D_{H}$) was obtained by first sampling a time series of $100,000$ time steps and computing dimension-wise Fast Fourier Transforms (using scipy.fft) for both the ground truth system and simulated time series, after discarding transients. The noise dominated high-frequency tails of the spectra were cut off, and the power spectra were slightly smoothed with a Gaussian kernel and normalized. We then computed the Hellinger distance (Mikhaeil et al., 2022) between smoothed power spectra of ground-truth, $F(\omega)$, and generated, $G(\omega)$, trajectories by
@@
To assess temporal agreement between model-generated and ground truth time series for ordinal and count observations, we computed a measure based on the average SACF. To this end we first sampled a time series of $100,000$ time steps and compute the dimension-wise Spearman autocorrelation for time lags up to $200$ for both generated and ground truth test data (see Fig. A9). The squared error between the resulting SACFs was then averaged across all dimensions.
@@
A mean squared prediction error (PE) was computed across test sets of length $T=10000$ by initializing the trained dendPLRNN with the test set time series up to some time point $t$ from the estimated intial state $\mathbf{E}[\mathbf{z}_{t}|\mathbf{x}_{1:t}]$, from where it was then iterated forward by $n$ time steps to yield a prediction at time step $t+n$. The $n$-step PE is then defined as the MSE between predicted and true observations:
@@
Due to exponential divergence of initially close trajectories in chaotic systems, the prediction error (PE) also depends on the precise initialization at time $t$. For SVAE training, an estimate of the initial state is directly provided by the encoder model. For BPTT-TF, the initial state $\mathbf{z}_{t}$ can be inferred by a jointly trained linear mapping from $\mathbf{z}_{t}$ to $\mathbf{x}_{t}$ (Brenner et al., 2022). For MTF and GVAE-TF, if $K=M$, the initial state $\mathbf{E}[\mathbf{z}_{t}|\mathbf{x}_{1:t}]$ is also directly given by the encoder model. If $K<M$, a subset of $M-K$ latent states is drawn from a standard normal distribution. In this under-specified case, we used a warm-up phase of $t_{w}$ time steps, where the system was initialized from the encoder at time $t-t_{w}$ and iterated forward, providing encoded states in the form of TF signals at every time step, which yields an initial state estimate $\mathbf{E}[\mathbf{z}_{t}|\mathbf{x}_{(t-t_{w}):t}]$. We determined $t_{w}=20$ by grid-search, leading to on average best results across datasets. We found that including this warm-up phase significantly improved prediction performance, although it still led to slightly worse results than for the fully specified case, $K=M$. Finally, for MS we neither have an encoder model nor a TF signal, and hence a sensible estimate for the initial state is difficult to obtain.
@@
To assess whether the global cross-correlation structure between the different ordinal time series is preserved by the reconstruction method, the Spearman correlation between each pair of ordinal time series was computed based on $100,000$ time steps long samples, using scipy.stats.spearmanr, for both generated and ground truth test data. The mean squared error between all elements of the correlation matrices was then taken.
@@
The maximum Lyapunov exponent quantifies the divergence rate of nearby trajectories, and (for discrete time systems $F_{\bm{\theta}}\big{(}{\bm{z}}_{t-1},{\bm{s}}_{t}\big{)}$ like the dendPLRNN, Eq.1) is defined as
@@
Figure A1: (a) Ground truth and rotated attractors of the Rössler system with associated $D_{stsp}$-values. (b) Correlation between geometrical reconstruction measures for the Rössler system directly in observation space given a co-trained linear (Gaussian) observation model ($D_{bin}$), and from a 3d PCA projection of latent space followed by an optimal rotation of the reconstructed attractor ($D_{PCA}$), based on a total of $30$ trained models.
@@ A.3 Comparison methods
As noted further above, different choices of DSR model are possible within our framework (see Suppl. Table B3). Since the focus here is on evaluating the training framework for DSR from multimodal data itself, for comparability we used the same dendPLRNN model in all five other methods compared to. The total number of trainable parameters was kept approximately the same for all methods (see also Table B1).
Sequential MVAE
First, we compared our method to a sequential MVAE as proposed in (Kramer et al., 2022), optimizing its multimodal ELBO. This, to our knowledge, is currently the only other general approach specifically designed for DSR from arbitrary data modalities observed simultaneously. We followed the implementation of the encoder model as provided on https://github.com/DurstewitzLab/mmPLRNN, training with a sequence length of $150$ time steps per batch, a hidden dimension of $20$, and other model parameters similar to the runs with MTF.
@@
Second, we tried ‘classical’ RNN training via truncated BPTT, where observations are provided as inputs at every time step, and the whole model is optimized using modality specific decoder models, eq. 19. Specifically, given observed multimodal time series $\bm{Y}$ of length $T$, observations $\bm{y}_{t}$ are included into the RNN updates according to
@@
where $\bm{U}$ is the input-to-hidden weight matrix. Training is performed on short subsequences of length $T_{seq}$, where $T_{seq}$ is a hyperparameter. The network then produces observations at each time step using the modality-specific decoder models, just as in MTF. During inference, the network is iterated forward by using the predicted observations at time $t$ as input for the predictions at time $t+1$. We probed different values for the sequence length $T_{seq}\in\{5\dots 50\}$. We found best settings to agree with the TF interval $\tau$ for MTF and the sequence length $L$ for multiple shooting, but – consistent with previous work (Mikhaeil et al., 2022) – plain error gradient/ sequence truncation often led to divergence and failed to capture the true time scales in the fMRI data.
@@
A third approach that can deal with multimodal observation models but, unlike TF, does not require model inversion, is ‘multiple shooting (MS)’ (Bock & Plitt, 1984), a method suggested in the dynamical systems literature (Voss et al., 2004). Multiple shooting aims to solve boundary value problems by dividing time into sub-intervals, treating each sub-interval as a separate initial value problem, and then imposing continuity conditions between intervals. Multiple shooting has also been applied to DS reconstruction and Neural-ODE training (Iakovlev et al., 2022), where the initial conditions (‘shooting nodes’) for each interval are model parameters, and continuity across intervals is enforced through a penalty term in the loss (Voss et al., 2004). Hence, rather than controlling trajectory flows through a sparse TF signal applied after forcing intervals $\tau$, alternatively one may reset the latent model trajectory to an inferred initial condition after $\tau$ time steps. The advantage is that this method does not require inversion of decoder models and is hence naturally suited to handle different data modalities without further care (i.e., retaining the distributional properties of the original data). More specifically, the observed time series $\bm{Y}$ is partitioned into $N_{seq}$ subsequences $\bm{Y}^{s},s=1\dots N_{seq}$, of length $L$, and for each subsequence a new initial condition $\bm{\mu}_{0}^{s}$ is learned. During training, trajectories are freely generated for $L$ time steps from $\bm{\mu}_{0}^{s}$ for each subsequence, and likelihoods for the observed trajectories $\bm{Y}^{s}$ are computed using the observation models from Eqs. 2-4. A consistency (penalty) term in the loss ensures continuity between subsequences according to
@@ A.4 Alternative encoder models
with means $\bm{\mu}_{\{g,o,c\}}$, diagonal covariances $\bm{\Sigma}_{\{g,o,c\}}$ and weights $w_{g}$ (Gaussian), $w_{o}$ (ordinal) and $w_{c}$ (count process), all set to $1/3$. We also tested a product-of-experts (PoE) approach (Hinton, 2002; Wu & Goodman, 2018), but found that it led to numerical instabilities for ordinal variables.
@@
The Transformer encoder is based on the architecture in (Vaswani et al., 2017). Since computation time scales with $T_{seq}^{2}$ in this architecture, we restricted the sequence length to $100$ steps. We used positional encodings to acknowledge the sequential (time series) nature of the data, as proposed in (Vaswani et al., 2017). The sequence plus embeddings is then passed through a standard Transformer encoder block using torch.nn.TransformerEncoder. As for the other encoder models, the output of the Transformer was mapped via linear readout layers onto the mean and logarithm of the covariance of the approximate posterior.
@@
We also tested a standard MLP with $3$ layers and rectified linear unit (ReLU) nonlinearity as encoder model, with outputs mapped via a linear readout layer to the mean and logarithm of the covariance of the approximate posterior, as for the other encoders tested.
@@ A.5 Datasets
Parameters used for producing ground truth data in the chaotic regime were $\sigma=10,\rho=28$, and $\beta=8/3$. Process noise was injected into the system by drawing from a Gaussian $\mathrm{d}\bm{\epsilon}\sim\mathcal{N}(\bm{0},0.01^{2}\mathrm{d}t\times\bm{I})$. For both training and test data, a trajectory of $100,000$ time steps was sampled, performing numerical integration with scipy.odeint ($\mathrm{d}t=0.05$; note this value differs from the one used in Mikhaeil et al. (2022), explaining the different $\tau$ values required). To obtain multimodal observations, trajectories drawn from the ground truth system were fed into the different types of observation models in Eqs. 2-4, with randomly drawn parameters.
@@
This same dataset was used previously in the study by Kramer et al. (2022) on multimodal data integration for DS reconstruction, and is openly available at https://github.com/DurstewitzLab/mmPLRNN. In the study, $26$ participants (of which $20$ were selected for analysis, cf. (Kramer et al., 2022)) were shown a series of images of different shapes (rectangles and triangles) while undergoing fMRI with a sampling rate of $1/3$ $\mathrm{Hz}$. The subjects were asked to identify the type of shape presented in the current image or in the preceding one, respectively, under three different task conditions: a continuous delayed response 1-back task (CDRT), a continuous matching 1-back task (CMT), and a 0-back control choice reaction task (CRT). A resting condition and an instruction phase were also included. Each task condition was repeated five times, where the last repetition of all task stages was left out as a test set in our analyses (see Fig. A2(b)). The brain activity of the participants as assessed via the BOLD signal, projected onto the first principal component within each of $10$ different brain regions in both hemispheres that were identified to be relevant for this the task. More details on this study can be found in Koppe et al. (2014). Due to the short length of these time series, for the fMRI results from Table 2 we chose a sequence length of $72$ time steps for testing and training.

### v3

@@ A.1 MTF framework
The MVAE consists of a set of decoder models, $p_{\bm{\theta}}(\bm{Y}|\bm{\tilde{Z}})$, shared with the DSR model, with examples given in Eqs. 6-8 and above, a prior model $p_{\bm{\theta}}(\bm{\tilde{Z}})$ instantiated by the DSR model, see Eq. 10, and the encoder (approximate posterior) $q_{\bm{\phi}}(\bm{\tilde{Z}}|\bm{Y})$, where a sample from this distribution provides the TF signal (see also Fig. 1). This setup is modular and all model components may be flexibly chosen. We assume conditional independence (given latent states) of the observations, such that the likelihood terms related to the observations simply sum up, i.e. $\log p_{\bm{\theta}}(\bm{{Y}}|\bm{\tilde{Z}})=\sum_{t=1}^{T}\left(\log p_{\bm{\theta}}(\bm{x}_{t}|\bm{\tilde{z}}_{t})+\log p_{\bm{\theta}}(\bm{o}_{t}|\bm{\tilde{z}}_{t})+\log p_{\bm{\theta}}(\bm{{c}_{t}}|\bm{\tilde{z}}_{t})\right)$ for the decoders specified in Eqs. 6-8 (with parameters shared with Eqs. 2-4).
For the encoder part, as the latent dendPLRNN process itself is conditionally Gaussian, we make a Gaussian assumption for the variational density $q_{\bm{\phi}}(\bm{\tilde{Z}}|\bm{Y})=\mathcal{N}(\bm{\mu}_{\bm{\phi}}(\bm{Y}),\bm{\Sigma}_{\bm{\phi}}(\bm{Y}))$ to approximate the true posterior $p(\bm{\tilde{Z}}|\bm{Y})$, where mean and covariance are functions of the data. We further use the common mean field approximation to factorize $q_{\bm{\phi}}(\bm{\tilde{Z}}|\bm{Y})$ across time (Girin et al., 2021). The mean $\bm{\mu}_{\theta}(\bm{Y})$ and covariance $\bm{\Sigma}_{\theta}(\bm{Y})$ of the approximate posterior are parameterized by one-dimensional (across time) convolutional layers (Brenner et al., 2022), using observations within the window defined by the input kernel. For the mean, we use a 4-layer stack of convolutional layers with decreasing kernel sizes $11,8,5,3$, while the diagonal covariance was parameterized by a single convolutional layer with a kernel size of $11$. We also tested various other choices for the encoder, as discussed in Appx. A.4.
@@
For the initial condition $\bm{z}_{1}$ of the latent process, if $K<M$, the $M-K$ remaining states are randomly sampled from a standard normal distribution. As shown in Mikhaeil et al. (2022), the best tradeoff between exploding gradients and capturing relevant long-term dependencies is achieved when choosing the forcing interval $\tau$ according to the system’s maximal Lyapunov exponent (predictability time). However, for non-continuous data (like counts), this cannot readily be determined, in which case the forcing interval $\tau$ may be regarded as a hyper-parameter determined via grid search (see also Appx. Fig. A7 for example reconstructions for different choices of $\tau$).
@@ A.2 Performance measures
The power spectrum Hellinger distance ($D_{H}$) was obtained by first sampling a time series of $100,000$ time steps and computing dimension-wise Fast Fourier Transforms (using scipy.fft) for both the ground truth system and simulated time series, after discarding transients. The noise-dominated high-frequency tails of the spectra were cut off, and the power spectra were slightly smoothed with a Gaussian kernel and normalized. We then computed the Hellinger distance (Mikhaeil et al., 2022) between smoothed power spectra of ground-truth, $F(\omega)$, and generated, $G(\omega)$, trajectories by
@@
To assess temporal agreement between model-generated and ground truth time series for ordinal and count observations, we computed a measure based on the average SACF, defined as:
$SACF_{i}(\tau)=\frac{\sum_{t=1}^{T-\tau}(r_{i,t}-\overline{r}_{i})\cdot(r_{i,t+\tau}-\overline{r}_{i})}{\sum_{t=1}^{T}(r_{i,t}-\overline{r}_{i})^{2}},$ (27)
For this, we first sampled a time series of $100,000$ time steps and computed the dimension-wise Spearman autocorrelation for time lags up to $200$ for both generated and ground truth test data (see Fig. A10). We then calculate the average squared error between the corresponding SACFs of ground truth and generated time series across all lags and dimensions as our performance metric:
$\text{MSE}_{SACF}=\frac{1}{NT}\sum_{i=1}^{N}\sum_{\tau=1}^{T}\left(\text{SACF}_{\text{gen},i}(\tau)-\text{SACF}_{\text{ground truth},i}(\tau)\right)^{2}$ (28)
@@
A mean squared prediction error (PE) was computed across test sets of length $T=10000$ by initializing the trained dendPLRNN with the test set time series up to some time point $t$ from the estimated initial state $\mathbf{E}[\mathbf{z}_{t}|\mathbf{x}_{1:t}]$, from where it was then iterated forward by $n$ time steps to yield a prediction at time step $t+n$. The $n$-step PE is then defined as the MSE between predicted and true observations:
@@
Given exponential trajectory divergence, the prediction error (PE) also depends on the precise initialization of the DS model at time $t$. For SVAE training, an estimate of the initial state is directly provided by the encoder model. For BPTT-TF, the initial state $\mathbf{z}_{t}$ can be inferred by a jointly trained linear mapping from $\mathbf{z}_{t}$ to $\mathbf{x}_{t}$ (Brenner et al., 2022). For MTF and GVAE-TF, if $K=M$, the initial state $\mathbf{E}[\mathbf{z}_{t}|\mathbf{x}_{1:t}]$ is also directly given by the encoder model. If $K<M$, a subset of $M-K$ latent states is drawn from a standard normal distribution. In this under-specified case, we used a warm-up phase of $t_{w}$ time steps, where the system was initialized from the encoder at time $t-t_{w}$ and iterated forward, providing encoded states in the form of TF signals at every time step, which yields an initial state estimate $\mathbf{E}[\mathbf{z}_{t}|\mathbf{x}_{(t-t_{w}):t}]$. We determined $t_{w}=20$ by grid-search, leading to on average best results across datasets. We found that including this warm-up phase significantly improved prediction performance, although it still led to slightly worse results than for the fully specified case, $K=M$. Finally, for MS we neither have an encoder model nor a TF signal, and hence a sensible estimate for the initial state is difficult to obtain.
@@
To assess whether the global cross-correlation structure between the different ordinal time series is preserved by the reconstruction method, the Spearman correlation between each pair of ordinal time series was computed based on $100,000$ time steps long samples, using scipy.stats.spearmanr, for both generated and ground truth test data.
$\text{SCC}_{ij}=\frac{1}{T}\sum_{t=1}^{T}\left(\frac{r_{it}-\overline{r}_{i}}{\sigma_{r_{i}}}\right)\left(\frac{r_{jt}-\overline{r}_{j}}{\sigma_{r_{j}}}\right)$ (31)
with $\sigma_{r_{i}}$ and $\sigma_{r_{j}}$ the standard deviation of the ranks of time series $i$ and $j$. We then calculated the MSE across all elements of the respective correlation matrices for the $N$ ordinal time series:
$\text{MSE}_{SCC}=\text{MSE}_{SCC}=\frac{1}{N^{2}-N}\sum_{i=1}^{N}\sum_{j\neq i}^{N}\left(\text{SCC}_{\text{gen},ij}-\text{SCC}_{\text{ground truth},ij}\right)^{2}$ (32)
@@
The maximum Lyapunov exponent quantifies the divergence rate of nearby trajectories, and (for discrete-time systems $F_{\bm{\theta}}\big{(}{\bm{z}}_{t-1},{\bm{s}}_{t}\big{)}$ like the dendPLRNN, Eq. 1) is defined as
@@
Figure A1: a: Ground truth and rotated attractors of the Rössler system with associated $D_{stsp}$-values. b: Correlation between geometrical reconstruction measures for the Rössler system directly in observation space given a co-trained linear (Gaussian) observation model ($D_{bin}$), and from a 3d PCA projection of latent space followed by an optimal rotation of the reconstructed attractor ($D_{PCA}$), based on a total of $30$ trained models.
@@ A.3 Comparison methods
As noted further above, different choices of DSR model are possible within our framework (see Appx. Table B3). Since the focus here is on evaluating the training framework for DSR from multimodal data itself, for comparability we used the same dendPLRNN model in all five other methods compared to. The total number of trainable parameters was kept approximately the same for all methods (see also Table B1).
Multimodal SVAE
First, we compared our method to a multimodal SVAE as proposed in (Kramer et al., 2022), optimizing its multimodal ELBO. This, to our knowledge, is currently the only other general approach specifically designed for DSR from arbitrary data modalities observed simultaneously. We followed the implementation of the encoder model as provided on https://github.com/DurstewitzLab/mmPLRNN, training with a sequence length of $150$ time steps per batch, a hidden dimension of $20$, and the parameters of the dendPLRNN the same as for the runs with MTF.
@@
Second, we tried ‘classical’ RNN training via truncated BPTT, where observations are provided as inputs at every time step, and the whole model is optimized using modality-specific decoder models, Eq. 22. Specifically, given observed multimodal time series $\bm{Y}$ of length $T$, observations $\bm{y}_{t}$ are included into the RNN updates according to
@@
where $\bm{U}$ is the input-to-hidden weight matrix. Training is performed on short subsequences of length $T_{seq}$, where $T_{seq}$ is a hyper-parameter. The network then produces observations at each time step using the modality-specific decoder models, just as in MTF. During inference, the network is iterated forward by using the predicted observations at time $t$ as input for the predictions at time $t+1$. We probed different values for the sequence length $T_{seq}\in\{5\dots 50\}$. We found best settings to agree with the TF interval $\tau$ for MTF and the sequence length $L$ for multiple shooting, but – consistent with previous work (Mikhaeil et al., 2022) – plain error gradient/ sequence truncation often led to divergence and failed to capture the true time scales in the fMRI data.
@@
A third approach that can deal with multimodal observation models but, unlike TF, does not require model inversion, is ‘multiple shooting (MS)’ (Bock & Plitt, 1984), a method suggested in the dynamical systems literature (Voss et al., 2004). Multiple shooting aims to solve boundary value problems by dividing time into sub-intervals, treating each sub-interval as a separate initial value problem, and then imposing continuity conditions between intervals. Multiple shooting has also been applied to DS reconstruction and Neural-ODE training (Iakovlev et al., 2022), where the initial conditions (‘shooting nodes’) for each interval are model parameters and continuity across intervals is enforced through a penalty term in the loss (Voss et al., 2004). Hence, rather than controlling trajectory flows through a sparse TF signal applied after forcing intervals $\tau$, alternatively one may reset the latent model trajectory to an inferred initial condition after $\tau$ time steps. The advantage is that this method does not require the inversion of decoder models and is hence naturally suited to handle different data modalities without further care (i.e., retaining the distributional properties of the original data). More specifically, the observed time series $\bm{Y}$ is partitioned into $N_{seq}$ subsequences $\bm{Y}^{s},s=1\dots N_{seq}$, of length $L$, and for each subsequence a new initial condition $\bm{\mu}_{0}^{s}$ is learned. During training, trajectories are freely generated for $L$ time steps from $\bm{\mu}_{0}^{s}$ for each subsequence, and likelihoods for the observed trajectories $\bm{Y}^{s}$ are computed using the observation models from Eqs. 2-4. A consistency (penalty) term in the loss ensures continuity between subsequences according to
@@ A.4 Alternative encoder models
with means $\bm{\mu}_{\{g,o,c\}}$, diagonal covariances $\bm{\Sigma}_{\{g,o,c\}}$ and weights $w_{g}$ (Gaussian), $w_{o}$ (ordinal) and $w_{c}$ (count process), all set to $1/3$. We also tested a product-of-experts (PoE) approach (Hinton, 2002; Wu & Goodman, 2018), where the estimates of the individual experts are multiplied instead of summed, which however often led to numerical instabilities during training on discrete variables.
@@
The Transformer encoder is based on the architecture in (Vaswani et al., 2017). Since computation time scales with $T_{seq}^{2}$ in this architecture, we restricted the sequence length to $100$ steps. We used positional encodings to acknowledge the sequential (time series) nature of the data, as proposed in (Vaswani et al., 2017). The sequence plus encodings is then passed through a standard Transformer encoder block using torch.nn.TransformerEncoder. As for the other encoder models, the output of the Transformer was mapped via linear readout layers onto the mean and logarithm of the covariance of the approximate posterior.
@@
We also tested a standard MLP with $3$ layers and rectified linear unit (ReLU) nonlinearity as an encoder model, with outputs mapped via a linear readout layer to the mean and logarithm of the covariance of the approximate posterior, as for the other encoders tested.
@@ A.5 Datasets
Parameters used for producing ground truth data in the chaotic regime were $\sigma=10,\rho=28$, and $\beta=8/3$. Process noise was injected into the system by drawing from a Gaussian $\mathrm{d}\bm{\epsilon}\sim\mathcal{N}(\bm{0},0.01^{2}\mathrm{d}t\times\bm{I})$. For both training and test data, a trajectory of $100,000$ time steps was sampled, performing numerical integration with scipy.odeint ($\mathrm{d}t=0.05$; note this value differs from the one used in Mikhaeil et al. (2022), explaining why different $\tau$ values were required for optimal reconstructions). To obtain multimodal observations, trajectories drawn from the ground truth system were fed into the different types of observation models in Eqs. 2-4, with randomly drawn parameters.
@@
This same dataset was used previously in the study by Kramer et al. (2022) on multimodal data integration for DS reconstruction, and is openly available at https://github.com/DurstewitzLab/mmPLRNN. In the study, $26$ participants (of which $20$ were selected for analysis, cf. Kramer et al. (2022)) were shown a series of images of different shapes (rectangles and triangles) while undergoing fMRI with a sampling rate of $1/3$ $\mathrm{Hz}$. The subjects were asked to identify the type of shape presented in the current image or in the preceding one, respectively, under three different task conditions: a continuous delayed response 1-back task (CDRT), a continuous matching 1-back task (CMT), and a 0-back control choice reaction task (CRT). A resting condition and an instruction phase were also included. Each task condition was repeated five times, where the last repetition of all task stages was left out as a test set in our analyses (see Fig. A2b). The brain activity of the participants as assessed via the BOLD signal projected onto the first principal component within each of $10$ different brain regions in both hemispheres that were identified to be relevant for this task. More details on this study can be found in Koppe et al. (2014). Due to the short length of these time series, for the fMRI results from Table 2 we chose a sequence length of $72$ time steps for testing and training.